In Part A, we saw that ΔG∘=−242.1 kJ for the hydrogenation of acetylene under standard conditions (all pressures equal to 1 atm and the common reference temperature 298 K). In Part B, you will determine the ΔG for the reaction under a given set of nonstandard conditions.

At 25 ∘C the reaction from Part A has a composition as shown in the table below. Substance Pressure (atm) C2H2(g) 3.75 H2(g) 4.25 C2H6(g) 5.25×10−2 What is the free energy change, ΔG, in kilojoules for the reaction under these conditions?
Answer #1

We find   Q of reaction , Q means reaction quotient

C2H2 (g) + 2H2 (g)----> C2H6 ( g )

Q   =   pC2H6 / ( pC2H2 x pH2^2)    = ( 0.0525 / 3.75 x 4.25^2)   = 0.0007750865

Now dG = dGo + RT ln K

dG = -242100 + ( 8.314 x 298) ln ( 0.0007750865)             ( dGo units taken in J as we have R in J)

= -242100-17745.7

    = -259846 J =   - 259.85 KJ